# Two-Player Zero-Sum Games with Bandit Feedback

### v3

@@ 5 ETC-TPZSG-AE Algorithm
The action pair $(i^{*},j^{*})$ is an $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE) if the following condition holds:
$A(i,j^{*})-\varepsilon\leq A(i^{*},j^{*})\leq A(i^{*},j)+\varepsilon,\quad\forall i\in S_{x},j\in S_{y}.$ (20)
@@
$k_{t}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}\Big\rceil$ (21)
@@
$\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}$ (22)
where $\varepsilon_{t}$ reflects how much players are allowed to deviate from the equilibrium and the tolerance over time gradually decreases to ensure convergence towards the NE. Substituting condition $k_{t}\geq\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}$ into the equation, where $k_{t}$ is defined in (21), we obtain $\varepsilon_{t}\leq\hat{\Delta}_{t}/2$. This bound indicates that as the number of rounds $t$ increases, the tolerance $\varepsilon_{t}$ becomes smaller, forcing the system to narrow the set of action pairs by progressively eliminating those that do not approximate a NE. As a result, over time, the algorithm keeps fewer action pairs, effectively focusing the exploration on action pairs that are close to the optimal one.
Since we assume that a unique pure NE exists in the game, a well estimated game matrix $\hat{A}$ should also exhibit the properties necessary to support the equilibrium. Specifically, $\hat{A}$ must satisfy the conditions required for the $\varepsilon$-NE in (20) to hold. Based on this principle, Algorithm 2 systematically eliminates action pairs that fail to satisfy this condition, as such pairs cannot be near to the equilibrium. By reducing the set of action pairs to include only the pairs that meet this criterion, it ensures that the remaining action pairs are the only ones that could potentially be near the NE; thus, it enables to focus on the most relevant action pairs. Consequently, we expect that the algorithm converges to an accurate approximation of the NE, even in the presence of estimation errors in $\hat{A}$.
@@
The Nash regret of Algorithm 2, interacting with $\sigma$-subgaussian payoffs, is upper bounded by
$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}\Big)+\max_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T$ (23)
@@
To analyze the regret of Algorithm 2, at first we define $t_{ij}=\min\{t:\hat{\Delta}_{t}<\Delta_{ij}/2\}$ as the earliest round satisfying $\hat{\Delta}_{t}<\Delta_{ij}/2$ and then we consider four distinct cases. First, we analyze the suboptimal action pairs which are not eliminated in round $t_{ij}$ or before. Next, we examine the suboptimal pairs that are eliminated in the round $t_{ij}$ or earlier, allowing us to bound the number of plays with suboptimal action pairs. The third case examines the scenario where the optimal action pair is eliminated in $t_{ij}$. Finally, we consider the case in which a suboptimal action pair remains as a unique one in the game (discussed in detail in Appendix C). ∎
@@
4:  if $\mathbf{|S_{t}|>1}:$ then $\triangleright$ Exploration Phase
5:   Explore each $(i,j)$ in $S_{t}$ $k_{t}$ times, defined in (21)
@@
10:  Remove the action pairs $(i,j)$ from $S_{t}$ NOT satisfying the following for $\varepsilon_{t}$ in (22): $\triangleright$ Elimination Phase
$\hskip 35.56593pt\hat{A}(i^{\prime},j)-\varepsilon_{t}\leq\hat{A}(i,j)\leq\hat{A}(i,j^{\prime})+\varepsilon_{t},\forall i^{\prime},\forall j^{\prime}$
@@
Theorem 3.
$R_{T}\leq\sum_{(i,j)\in S_{A_{1}}}\Big(\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{256\sigma^{2}}{\Delta_{ij}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\frac{512\sigma^{2}}{\lambda}\Big)+\lambda T.$ (24)
Proof.
Although we have a different notion of regret, this does not affect the underlying structure of the regret analysis. This ensures that the previous theoretical bounds can be extended using $\Delta_{ij}^{*}\leq\Delta_{ij}$, then the result follows. ∎
We simplify the regret bound to $O(\frac{\log(T\Delta^{2})}{\Delta})$, as the logarithmic term dominates since by setting $\lambda=\sqrt{16\sigma^{2}e/T}$, we ensure that the term $\lambda T$ is bounded by $\sqrt{16\sigma^{2}eT}$, which is at most $\frac{e}{\Delta_{ij}}$ when $\Delta_{ij}\leq\lambda$. Our results are consistent with the regret bound derived by [6].
Compared to our regret bound $O(\frac{\log(T\Delta^{2})}{\Delta})$ with the bound $O(\frac{\log_{+}(T\Delta^{2})}{\Delta})$ in [18], although they seem very similar, we observe that our study introduces a more refined notion of the suboptimality gap $\Delta$. Specifically, in [18], $\Delta$ refers to the gap between pure NE and the action chosen when the opponent plays the action in equilibrium. Moreover, our regret bound is tighter, as their formulation involves the square of the sum of suboptimality gaps, while ours uses the sum of squares of suboptimality gaps within the logarithmic factor. Moreover, our elimination algorithm specifically incorporates the extended pure NE property through $\varepsilon$, which also allows us to explain why we achieve a tighter regret bound for a TPZSG setting where both players have a pure strategy.
As a result, we have $R_{T}^{*}\leq R_{T}$, where $R_{T}^{*}$ evaluates its performance against the NE, capturing the cumulative loss relative to the equilibrium strategy, while the regret analysis of $R_{T}$ allows us to consider the losses of both players with respect to their best individual responses in each round, providing a broader and more flexible notion of regret. This outcome aligns with the interpretation that the expected regret against the NE establishes a more specific comparison, and thus, it naturally provides a tighter upper bound on the expected regret.
@@ 6 ETC-TPZSG-AE-NUE Algorithm
As mentioned in Section 3, the strategies of the row and column players are defined as probability distributions $p$ and $q$, respectively, over their corresponding action sets. In round $t$, using the estimated game matrix, we denote the estimated strategy of the row player by $\hat{p}_{t}$ and the estimated strategy of the column player by $\hat{q}_{t}$. Since the true game matrix is assumed to have the pure NE, we expect these estimated strategies to converge to the pure strategies over time.
The estimated strategies of the players serve as a basis to determine the exploration time in Algorithm 3. The probability of selecting an action pair $(i,j)$ is estimated by $\hat{P}_{ij,t}=\hat{p}^{*}_{i,t}\hat{q}^{*}_{j,t}$, which satisfies $\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=1$, where $(\hat{p}_{t}^{*},\hat{q}_{t}^{*})$ is a mixed NE in round $t$. In other words, once optimal strategies are determined, we calculate the joint probability $\hat{P}_{ij,t}$, which represents the probability that the row player chooses action $i$ and the column player chooses action $j$ in round $t$. Consequently, in Algorithm 3, the exploration time for each action pair $(i,j)$ is proportional to $\hat{P}_{ij,t}$, allowing action pairs with higher estimated probabilities to be played more frequently.
We incorporate the strategies of the players $\hat{p}^{*}_{t}$ and $\hat{q}^{*}_{t}$ together with $k_{t}$. Specifically, in each round $t$, this involves estimating the optimal strategies for each player by solving a linear system using the estimated payoff matrix $\hat{A}$ as follows:
$\begin{aligned} &\\ &\hat{P}_{ij,t}=\hat{p}^{*}_{i,t}\hat{q}^{*}_{j,t}\\ &\hat{p}^{*}_{t}=\operatorname*{argmax}_{\hat{p}}\min_{\hat{q}}\hat{p}^{T}\hat{A}\hat{q}\\ &\hat{q}^{*}_{t}=\operatorname*{argmin}_{\hat{q}}\max_{\hat{p}}\hat{p}^{T}\hat{A}\hat{q}\\ \end{aligned}\qquad\begin{aligned} &\text{Subject to:}\\ &\sum_{i\in S_{x}}\hat{p}_{i,t}=1,\quad\hat{p}_{i,t}\geq 0,\ \forall i\in S_{x}\\ &\sum_{j\in S_{y}}\hat{q}_{j,t}=1,\quad\hat{q}_{j,t}\geq 0,\ \forall j\in S_{y}\\ &\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=\sum_{i,j}\hat{p}_{i,t}\hat{q}_{j,t}=1\end{aligned}$ (25)
Thus, the exploration time is given by the following:
$k_{ij,t}=k_{t}+\Big\lceil\frac{16\sigma^{2}\hat{P}_{ij,t}}{\hat{\Delta}_{t}^{2}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}\Big\rceil$ (26)
where $k_{t}$ is defined in (21). Thus, the algorithm explores each action pair at least $k_{t}$ times. Since players’ strategies are expected to converge to pure ones over time, the Algorithm 3 progressively focuses on playing the optimal action pair more. This ensures that action pairs with higher estimated probabilities are explored more frequently, while those with lower probabilities are selected less often. Thus, these algorithms achieve a more efficient balance between exploration and exploitation by reducing unnecessary plays on suboptimal action pairs, which accelerates convergence toward optimal action pair.
@@
4:  if $\mathbf{|S_{t}|>1}:$ then $\triangleright$ Exploration Phase
5:   Explore each $(i,j)$ in $S_{t}$ $k_{ij,t}$ times, defined in (26)
6:   Update $\hat{A}$ using (4)
[…]
8:   Play with a unique action pair in $S_{t}$ until step $T$
9:  end if
10:  Remove the action pairs $(i,j)$ from $S_{t}$ NOT satisfying the following for $\varepsilon_{t}$ in (22): $\triangleright$ Elimination Phase
$\hskip 35.56593pt\hat{A}(i^{\prime},j)-\varepsilon_{t}\leq\hat{A}(i,j)\leq\hat{A}(i,j^{\prime})+\varepsilon_{t},\forall i^{\prime},\forall j^{\prime}$
11:  Solve the linear system in (25) to get $\hat{P}_{ij,t+1}$
12:  $S_{t+1}\leftarrow S_{t}$
13:  $\hat{\Delta}_{t+1}\leftarrow\frac{\hat{\Delta}_{t}}{2}$ where $\hat{\Delta}_{t}=2^{-t+2}\sigma$
14:end for
@@
The Nash regret of Algorithm 3 is upper bounded by
$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}\Big)+\max_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T$ (27)
@@ 8 Discussion
We investigate a TPZSG setting with bandit feedback, where the true payoff matrix is unknown and must be learned through player interactions. This setting is challenging, as players must estimate payoffs while making strategic decisions in an adversarial environment. We adopt the ETC algorithm due to its simplicity, widespread use and lack of prior analysis in this context. We integrate it to an adaptive elimination method, allowing the systematic removal of suboptimal action pairs, thereby improving convergence to the equilibrium and reducing unnecessary exploration on suboptimal ones. Finally, we combine the adaptive elimination algorithm with a non-uniform exploration approach to explore action pairs more effectively. A key contribution of our work is the derivation of instance-dependent upper bounds on the expected regret for proposed algorithms, which has received limited attention in the literature on zero-sum games.
@@ Appendix A Useful Properties
$\mathbb{P}(X\geq\epsilon)\leq\exp{\Big(-\frac{\epsilon^{2}}{2\sigma^{2}}\Big)}.$ (31)
@@ Appendix B Proof of Theorem 1
$\leq\sqrt{\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}}{4\sigma^{2}}\Big)}\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{4\sigma^{2}}\Big)}}$ (37)
$=\sqrt{\exp{\Big(-\frac{k[(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}]}{4\sigma^{2}}\Big)}}$ (38)
$\leq\sqrt{\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{8\sigma^{2}}\Big)}}$ (39)
$=\sqrt{\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}})^{2}}{8\sigma^{2}}\Big)}}$ (40)
$=\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}})^{2}}{16\sigma^{2}}\Big)}$ (41)
@@
To obtain (33), we utilize the fact that the probability of the intersection of multiple events is at most that of any individual event, as stated in 3. For step (37), we apply Lemma 4 with $\sqrt{2\sigma^{2}/k}$-subgaussian random variable. At this point, we need to show that the differences $\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})-\hat{A}(i_{1},j^{\prime})+A(i_{1},j^{\prime})$ and $\hat{A}(i^{\prime},j_{1})-A(i^{\prime},j_{1})-\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})$ are $\sqrt{2\sigma^{2}/k}$-subgaussian. To establish this, we utilize Lemma 5, which provides key properties of subgaussian random variables. Specifically, for any $(i,j)\in S_{A}$, $\hat{A}(i,j)-A(i,j)$ are $\sqrt{\sigma^{2}/k}$-subgaussian since during the exploration phase, each action pair is played $k$ times. Then by Lemma 5, the difference of two independent subgaussian random variables with parameters $\sqrt{\sigma^{2}/k}$ is subgaussian with parameter $\sqrt{2\sigma^{2}/k}$. To obtain the last inequality (39), we apply Lemma 2.
@@
$\leq\min\Big\{\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}}{4\sigma^{2}}\Big)},\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{4\sigma^{2}}\Big)}\Big\}$ (45)
$=\exp{\Big(-\frac{k\max\{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2},(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\}}{4\sigma^{2}}\Big)}$ (46)
$\leq\exp{\Big(-\frac{k[(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}]}{8\sigma^{2}}\Big)}$ (47)
$\leq\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{16\sigma^{2}}\Big)}$ (48)
$\leq\exp{\Big(-\frac{k\Delta_{i^{\prime}j^{\prime}}^{2}}{16\sigma^{2}}\Big)}$ (49)
@@
$\mathbb{E}[n_{ij,T}]\leq k+(T-Nk)\exp{\Big(-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\Big)}.$ (51)

### v4

@@ 5 ETC-TPZSG-AE Algorithm
The action pair $(i^{*},j^{*})$ is an $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE) if the following condition holds for any $\varepsilon>0$ and for all $i\in S_{x}$ and $j\in S_{y}$:
$A(i,j^{*})-\varepsilon\leq A(i^{*},j^{*})\leq A(i^{*},j)+\varepsilon.$ (17)
@@
$k_{t}=\left\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}}\ln{\left(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\right)}\right\rceil$ (18)
@@
$\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{\left(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\right)}}$ (19)
where $\varepsilon_{t}$ reflects how much players are allowed to deviate from the equilibrium and the tolerance over time gradually decreases to ensure convergence towards the NE. Substituting condition $k_{t}\geq\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}}\ln{\left(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\right)}$ into the equation, where $k_{t}$ is defined in (18), we obtain $\varepsilon_{t}\leq\hat{\Delta}_{t}/2$. This bound indicates that as the number of rounds $t$ increases, the tolerance $\varepsilon_{t}$ becomes smaller, forcing the system to narrow the set of action pairs by progressively eliminating those that do not approximate a NE. As a result, over time, the algorithm keeps fewer action pairs, effectively focusing the exploration on action pairs that are close to the optimal one.
Since we assume that a unique pure NE exists in the game, a well estimated game matrix $\hat{A}$ should also exhibit the properties necessary to support the equilibrium. Specifically, $\hat{A}$ must satisfy the conditions required for the $\varepsilon$-NE in (17) to hold. Based on this principle, Algorithm 2 systematically eliminates action pairs that fail to satisfy this condition, as such pairs cannot be near to the equilibrium. By reducing the set of action pairs to include only the pairs that meet this criterion, it ensures that the remaining action pairs are the only ones that could potentially be near the NE; thus, it enables to focus on the most relevant action pairs. Consequently, we expect that the algorithm converges to an accurate approximation of the NE.
@@
For $1\leq ml\ll T$, the Nash regret of Algorithm 2, interacting with $\sigma$-subgaussian payoffs, is upper bounded by
$R_{T}^{*}\leq$ $\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\left(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right)+\sum_{(i,j)\in S_{A_{2}}}\left(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}\right)+\max_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T$ (20)
@@
We first define $t_{ij}=\min\{t:\hat{\Delta}_{t}<\Delta_{ij}/2\}$ as the earliest round satisfying $\hat{\Delta}_{t}<\Delta_{ij}/2$. Thus, we consider four distinct cases. First, we analyze the suboptimal action pairs which are not eliminated in round $t_{ij}$ or before. Next, we examine the suboptimal pairs that are eliminated in the round $t_{ij}$ or earlier, allowing us to bound the number of plays with suboptimal action pairs. The third case examines the scenario where the optimal action pair is eliminated in $t_{ij}$. Finally, we consider the case in which a suboptimal action pair remains as a unique one in the game. (See Appendix C) ∎
Theorem 3.
$R_{T}\leq$ $\sum_{(i,j)\in S_{A_{1}}}\left(\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{256\sigma^{2}}{\Delta_{ij}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right)+\sum_{(i,j)\in S_{A_{2}}}\left(\frac{512\sigma^{2}}{\lambda}\right)+\lambda T.$ (21)
Proof.
Although the regret notion differs, this does not affect the underlying regret analysis. Thus, the previous theoretical bounds can be extended using Lemma 1, which completes the proof. ∎
@@
4:  if $\mathbf{|S_{t}|>1}:$ then
5:   Explore each $(i,j)$ in $S_{t}$ $k_{t}$ times in (18)
@@
10:  Remove all $(i,j)$ from $S_{t}$ NOT satisfying:
[…]
@@
We simplify the regret bound to $O\left(\frac{\log(T\Delta^{2})}{\Delta}\right)$, as the logarithmic term dominates since by setting $\lambda=\sqrt{16\sigma^{2}e/T}$, we ensure that the term $\lambda T$ is bounded by $\sqrt{16\sigma^{2}eT}$, which is at most $\frac{e}{\Delta_{ij}}$ when $\Delta_{ij}\leq\lambda$. Our results are consistent with the regret bound derived by [6].
Compared to our result, [18] achieves a regret bound of $O\left(\left(\sum_{i\neq i^{*}}\frac{1}{\Delta_{i}}+\sum_{j\neq j^{*}}\frac{1}{\Delta^{\prime}_{j}}\right)\log T\right)$ where $\Delta_{i}=A(i^{*},j^{*})-A(i,j^{*})$ and $\Delta^{\prime}_{j}=A(i^{*},j)-A(i^{*},j^{*})$ denote the gaps between pure NE and the payoff obtained by deviating while the opponent plays its equilibrium action. Although the bound is not directly comparable to ours, as the settings and the definition of suboptimality gaps differ, it currently presents the only instance-dependent regret analysis for general $m\times l$ zero-sum matrix games. Our regret bound is tighter and our analysis relies on a more refined notion of the suboptimality gap $\Delta$, as it exhibits a sharper dependence on the gap parameter and avoids the cumulative sum of suboptimality gaps that appears in [18]. Furthermore, our elimination algorithm specifically incorporates the extended pure NE property through $\varepsilon$ in a centralized setting, which explains why we achieve a tighter regret bound for a TPZSG setting where both players have a pure strategy; this is further supported by the simulation results in Section 7.
As a result, we have $R_{T}^{*}\leq R_{T}$, where $R_{T}^{*}$ evaluates its performance against the NE, capturing the cumulative loss relative to the equilibrium strategy, while the regret analysis of $R_{T}$ allows us to consider the losses of both players with respect to their best individual responses in each round, providing a broader and more flexible notion of regret. This outcome aligns with the interpretation that the expected regret against the NE establishes a more specific comparison, and thus, it naturally provides a tighter upper bound.
@@ 6 ETC-TPZSG-AE-NUE Algorithm
In round $t$, we denote the estimated optimal strategy of the row player by $\hat{p}_{t}$ and the estimated optimal strategy of the column player by $\hat{q}_{t}$, where they are calculated trough the estimated game matrix $\hat{A}$. We expect these estimated strategies to converge to the pure strategies over time.
The estimated strategies of the players serve as a basis to determine the exploration time in Algorithm 3. The joint probability of selecting an action pair $(i,j)$ is estimated by $\hat{P}_{ij,t}=\hat{p}_{i,t}\hat{q}_{j,t}$, which satisfies $\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=1$, where $(\hat{p}_{t},\hat{q}_{t})$ is a mixed NE in round $t$. Consequently, in Algorithm 3, the exploration time for each action pair $(i,j)$ is proportional to $\hat{P}_{ij,t}$, allowing action pairs with higher estimated probabilities to be played more frequently.
We incorporate the strategies of the players $\hat{p}_{t}$ and $\hat{q}_{t}$ together with $k_{t}$ in (18). Specifically, in each round $t$, this involves estimating the optimal strategies for each player by solving a linear system using the estimated payoff matrix $\hat{A}$. Thus, the exploration time is given by the following:
$k_{ij,t}=k_{t}+\left\lceil\frac{16\sigma^{2}\hat{P}_{ij,t}}{\hat{\Delta}_{t}^{2}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}\right\rceil$ (22)
Thus, the algorithm explores each action pair at least $k_{t}$ times. Since players’ strategies are expected to converge to pure ones over time, the Algorithm 3 progressively focuses on playing the optimal action pair more. This ensures that action pairs with higher estimated probabilities are explored more frequently, while those with lower probabilities are selected less often. Thus, these algorithms achieve a more efficient balance between exploration and exploitation by reducing unnecessary plays on suboptimal action pairs, which accelerates convergence toward optimal action pair.
@@
3:for $t=0,1,2,...,\left\lfloor\frac{1}{4}\log_{2}\frac{T}{e}\right\rfloor$ do
4:  Run Algorithm 2 with exploration time $k_{ij,t}$ in (22)
5:  Calculate optimal strategies $\hat{p}_{t}$ and $\hat{q}_{t}$ based on $\hat{A}$
6:  Update $\hat{P}_{ij,t+1}$
7:end for
@@
For $1\leq ml\ll T$, the Nash regret of Algorithm 3 is upper bounded by
$R_{T}^{*}\leq$ $\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\left(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right)+\sum_{(i,j)\in S_{A_{2}}}\left(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}\right)+\max_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T$ (23)
@@ 8 Discussion
We investigate a TPZSG setting with bandit feedback, where the true payoff matrix is unknown and must be learned through player interactions. This setting is challenging, as players must estimate payoffs while making strategic decisions with strictly opposing objectives. We adopt the ETC algorithm due to its simplicity, widespread use and lack of prior analysis in this context. We integrate it to an adaptive elimination method, allowing the systematic removal of suboptimal action pairs, thereby improving convergence to the equilibrium and reducing unnecessary exploration on suboptimal ones. Finally, we combine the adaptive elimination algorithm with a non-uniform exploration approach to explore action pairs more effectively. A key contribution of our work is the derivation of instance-dependent upper bounds on the expected regret for proposed algorithms, which has received limited attention in the literature on zero-sum games.
@@ Appendix A Useful Properties
$\mathbb{P}(X\geq\epsilon)\leq\exp{\left(-\frac{\epsilon^{2}}{2\sigma^{2}}\right)}.$ (27)
@@ Appendix B Proof of Theorem 1
$\leq\sqrt{\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}}{4\sigma^{2}}\right)}\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{4\sigma^{2}}\right)}}$ (33)
$=\sqrt{\exp{\left(-\frac{k[(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}]}{4\sigma^{2}}\right)}}$ (34)
$\leq\sqrt{\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{8\sigma^{2}}\right)}}$ (35)
$=\sqrt{\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}})^{2}}{8\sigma^{2}}\right)}}$ (36)
$=\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}})^{2}}{16\sigma^{2}}\right)}$ (37)
@@
To obtain (29), we utilize the fact that the probability of the intersection of multiple events is at most that of any individual event, as stated in 4. For step (33), we apply Lemma 5 with $\sqrt{2\sigma^{2}/k}$-subgaussian random variable. At this point, we need to show that the differences $\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})-\hat{A}(i_{1},j^{\prime})+A(i_{1},j^{\prime})$ and $\hat{A}(i^{\prime},j_{1})-A(i^{\prime},j_{1})-\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})$ are $\sqrt{2\sigma^{2}/k}$-subgaussian. To establish this, we utilize Lemma 6, which provides key properties of subgaussian random variables. Specifically, for any $(i,j)\in S_{A}$, $\hat{A}(i,j)-A(i,j)$ are $\sqrt{\sigma^{2}/k}$-subgaussian since during the exploration phase, each action pair is played $k$ times. Then by Lemma 6, the difference of two independent subgaussian random variables with parameters $\sqrt{\sigma^{2}/k}$ is subgaussian with parameter $\sqrt{2\sigma^{2}/k}$. To obtain the last inequality (35), we apply Lemma 3.
@@
$\leq\min\left\{\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}}{4\sigma^{2}}\right)},\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{4\sigma^{2}}\right)}\right\}$ (41)
$=\exp{\left(-\frac{k\max\{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2},(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\}}{4\sigma^{2}}\right)}$ (42)
$\leq\exp{\left(-\frac{k[(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}]}{8\sigma^{2}}\right)}$ (43)
$\leq\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{16\sigma^{2}}\right)}$ (44)
$\leq\exp{\left(-\frac{k\Delta_{i^{\prime}j^{\prime}}^{2}}{16\sigma^{2}}\right)}$ (45)
@@
$\mathbb{E}[n_{ij,T}]\leq k+(T-Nk)\exp{\left(-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\right)}.$ (47)